OpenStudy (anonymous):

solve with work please! 6y^2-24y=0

OpenStudy (anonymous):

6y(y-4)=0, 6y=0 or y-4=0, y=0 or y=4

OpenStudy (cwrw238):

facrtorise: 6y(y - 4)= 0 either 6y = 0 or y - 4 = 0 y = 0 or y = 4 solution set is (0,4)

OpenStudy (anonymous):

how do you get y-4=0 ? and 6y?? im so lost..

OpenStudy (anonymous):

what is the first step when solving this? :(